p(-15/17,-8/17) is found on the unit circle. Find sin 0 ( a line through the middle of 0)and cos 0 ( a line through the middle of 0)
Answer by Alan3354(21605)

p(-15/17,-8/17) is found on the unit circle. Find sin 0 ( a line through the middle of 0)and cos 0 ( a line through the middle of 0)
-------------
By "a line through the middle of O" I'll assume you mean from the point to the Origin.
Draw a unit circle about the origin and plot the point on it.
The sine of the angle (between the x-axis and the line) is the side opposite,
-8/17, over 1, and the cosine of the angle is the side adjacent over 1, = 15/17.
So sin O = -8/17 and the cos(O) = 15/17
Angle O = -28 degs, or 332 degrees.